WAC Notes July 20 2019

From PREX Wiki
Jump to: navigation, search
Documentation
HOW TOs for shift crew
Expert Tools
All Expert Contacts



PREX Main << Weekly Analysis Coordinator << WAC Notes << Instructions

WAC Notes July 19 2019 << >> WAC Notes July 21 2019

July 20 2019

Removed S0 -> begins Slug 8 at run 3453 (two arms) - 3454 (R arm trips) - 3486-> good 50 uA Left arm data

Production Run List - Swing 2019/07/19 - OWL 2019/07/20

  • Calibration - FFB Test - 3486: Large Aq, needs cuts probably on noisy BPM data
  • Calibration - FFB Test - 3487: Large Aq
  • Production - 3488
  • Production - 3489: After cuts no visible plot problems
  • Production - 3490
  • Production - 3491
  • Production - 3492
  • Production - 3493
  • Production - 3494: Mostly beam trip
  • Production - 3495
  • Production - 3496: Has broad BCM noise at the beginning of the run
  • Production - 3497: Has broad BCM noise in the middle of the run
  • Production - 3498
  • Production - 3499: Has broad BCM noise at end of run - it looks like this is a brief and transient beam current noise issue
  • Production - 3500
  • Production - 3501
  • Production - 3502
  • Production - 3503
  • Production - 3504
  • Production - 3505
  • Production - 3506: 2ppm Aq
  • Production - 3507: very short run, ends in beam off
  • No beam, junk - 3508->

I would like to arbitrarily start Slug 9 at 3520 (beam went away during 3508, 3509, and until )

Non-Production Runs and Bad Runs

Pedestal/Calibrations Updates

  • Pedestals for 70 uA running (Slug 5, runs 3424 - 3432) have been recomputed and those runs are respun now (3:00 PM)
    • Main det, BCMs (beamline), and SAMs all with that run range now.
    • Git rm Parity/prminput/prexCH_beamline_pedestal.3425-.map
  • Run 3520-> starts slug 9 and has all new pedestals and even changes bcm an ds 3 for ds 1 as the normalizer -> The Pedestals have been updated from 3707321 only for bcm an ds 1x
  • Fully removed the BPM 4e 4a and 12 stability cuts
  • Cleaning up BPM noise with stability cuts on X and Y HALOG

UPDATE

  • Changed all 3520 -> 3453 to respin slug 8 (50 uA) too

Issues

  • Slug 3 -> all have the wrong SAM pedestals
  • We need better Slug averaged plots, and Ciprian and Sakib are working on it
  • We have obtained plots for Slugs 1-8 using the list of parameters as it works now (~/PREX/prompt/collector/drawPostpan/drawPostPan.C(filename,descriptor,deviceList))
  • Right arm has been dead since run 3454 (2nd run of slug 8, and it was a needs-cuts run anyway)
  • FFB test during run 3486 and 3487
  • Runs 3489 and 3493 have some questionable regression during 1 minirun: too little data/noisy beam/FFB problems?? BPM 12 changes response after 3492
  • Runs 3496, 7, and 9 all have large RMS BCM noise
  • Runs 3498-3500 all have BPM4eX noise larger than normal
  • All slug 8 needed to be respun as an errant BPM 4a and 4e stability cut had been implemented, we are now needing to implement saturation cuts independently of stability cut for these BPMs (and maybe for 12 as well??)

Other Status Updates